# Boosting HDR Image Reconstruction via Semantic Knowledge Transfer

### v1

@@ 4.3 Qualitative Evaluation
PSNR-$\mu$uparrow SSIM-$\mu$uparrow PSNR-$\mu$uparrow SSIM-$\mu$uparrow
AHDR 41.92 0.9836 - -
w/o $\mathcal{L}_{\text{content}}$ 42.15 0.9841 42.69 0.9849
w/o $\mathcal{L}_{\text{color}}$ 43.09 0.9861 43.19 0.9864
w/o SKAM 43.11 0.9862 43.13 0.9873
w/o FastSAM 41.86 0.9829 41.86 0.9830
w HRNet 42.95 0.9859 42.95 0.9859
AHDR+ 43.18 0.9874 43.18 0.9874
Table 3: Ablation studies of proposed framework.
4.4 Ablation Studies
@@
Considering that Tel’s dataset offers a more balanced distribution of scenes and motions, We conduct ablation experiments on this dataset to validate the effectiveness of our proposed method. From Tab. 3, we can observe that the absence of any component in the semantic knowledge transfer scheme leads to a noticeable drop in performance. Specifically, $\mathcal{L}_{\text{content}}$ as the foundational loss function for self-distillation, has a significant impact on model performance. In other words, the proposed $\mathcal{L}_{\text{color}}$ effectively enhances the self-distillation capability when built upon the base loss function. W/o SKAM refers to the variant where the semantic knowledge alignment module is removed. We observe that this variation slightly degrades the performance of ORM, indicating that SKAM contributes to effective semantic knowledge transfer. Furthermore, w/o FastSAM indicates the removal of semantic priors in the SPGRM, instead employing a simple second-stage enhancement. We discover that merely increasing the model’s parameter count does not lead to substantial performance gains and may even cause overfitting, resulting in performance degradation. In addition, we assess the influence of diverse semantic priors derived from FastSAM and the widely utilized HRNet [30]. Given that SAM-based method offers richer and more nuanced semantic information across multiple scales, our approach outperforms HRNet that relies solely on semantic cues from instance and category labels.
5 Conclusion
@@
Existing multi-exposure HDR imaging methods primarily focus on mitigating artifact issues caused by motion, but they overlook the inherently ill-posed nature of captured LDR images in real-world scenarios. To tackle these challenges, we introduces a general framework that leverages semantic priors from pretrained segmentation models to boost existing HDR reconstruction models without modifying the original network. By endowing models with semantic knowledge, our framework enables a deeper understanding of image content, providing explicit guidance for the reconstruction process. Extensive experiments demonstrate that the proposed framework consistently improves the performance of multiple baseline models, achieving state-of-the-art results across diverse datasets encompassing various scenes and formats.

### v2

@@ IV-C Qualitative Evaluation
PSNR-PU $\uparrow$ SSIM-PU $\uparrow$ PSNR-PU $\uparrow$ SSIM-PU $\uparrow$
AHDR[49] 42.29 0.9919 - -
+ Feature Injection 42.45 0.9921 - -
+ Segmentation Map 42.39 0.9923 - -
w/o $\mathcal{L}_{\text{content}}$ 42.46 0.9920 42.98 0.9925
w/o $\mathcal{L}_{\text{color}}$ 43.41 0.9926 43.59 0.9928
w/o SKAM 43.53 0.9928 43.57 0.9931
Direct Alignment 39.69 0.9890 41.63 0.9907
w/o FastSAM[64] 42.23 0.9915 42.21 0.9913
w HRNet[41] 43.29 0.9929 43.28 0.9929
AHDR+ 43.61 0.9937 43.62 0.9937
IV-D Ablation Studies
Considering that Tel’s dataset offers a more balanced distribution of scenes and motions, We conduct ablation experiments on this dataset to validate the effectiveness of our proposed method.
Effects of Semantic Knowledge Transfer Scheme. To bridge the performance gap between the teacher and the student model, we introduce two loss functions: $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{content}}$. These losses facilitate semantic knowledge transfer at the color and content levels, respectively. To assess their impact on model performance, we conduct experiments by selectively removing these loss functions and analyzing the results. As presented in Tab IV, $\mathcal{L}_{\text{content}}$, which is a widely adopted approach in existing distillation schemes, serves as the foundational loss function for self-distillation. It significantly influences the model’s overall performance. Meanwhile, while relying solely on $\mathcal{L}_{\text{color}}$ does not suffice for efficient self-distillation, combining it with $\mathcal{L}_{\text{content}}$ yields a noticeable performance boost. This suggests that $\mathcal{L}_{\text{color}}$ effectively addresses the challenge of dramatic brightness variations across different regions of HDR images, enhancing the model’s ability to capture color consistency and detail in challenging lighting conditions.
Effects of Semantic Knowledge Alignment Module. In addition to model-level distillation, the Semantic Knowledge Alignment Module further enhances self-distillation by mapping features from different domains into a shared latent space for semantic alignment. As shown in Tab. IV, the variant labeled “W/o SKAM” refers to the model with the SKAM removed. We observe that this variation slightly degrades the performance of the ORM, suggesting that SKAM plays a key role in facilitating effective semantic knowledge transfer. Furthermore, we assessed the impact of directly aligning the features of different models using Mean Squared Error loss without SKAM, denoted as “Direct Alignment.” The results indicate that, due to the domain gap between the intermediate features of different models, direct alignment significantly impairs the performance of both the student and teacher models. In contrast, SKAM employs an encoder to map $F_{S}^{k}$ and $F_{T}^{k}$ to a shared latent space. A small loss in this setup demonstrates that $F_{S}^{k}$ can effectively reconstruct the masked parts of $F_{T}^{k}$ within the latent space, thereby achieving semantic alignment between $F_{T}^{k}$ and $F_{S}^{k}$.
Figure 7: Qualitative generalization results on Sen et al.’s dataset[36] and Tursun et al.’s dataset[40].
@@ IV-D Ablation Studies
Effects of Semantic Knowledge Bank. The proposed framework aims to address the ill-posed nature of High Dynamic Range reconstruction by leveraging semantic knowledge from the Semantic Knowledge Bank. To validate its effectiveness, we performed ablation studies by removing the semantic priors and extracting the semantic priors from different models. As shown in Tab.IV, “w/o FastSAM” indicates the removal of semantic priors in the SPGRM, instead employing a simple second-stage enhancement strategy. The results demonstrate that merely increasing the model’s parameter count does not lead to substantial performance gains and may even cause overfitting, resulting in performance degradation. In addition, we evaluated the influence of various semantic priors derived from FastSAM and the widely used HRNet [41]. Given that the SAM-based method provides richer and more nuanced semantic information across multiple scales, our approach outperforms HRNet, which relies solely on semantic cues from instance and category labels.
Effects of Semantic Prior Injection Strategy. We conducted additional ablation experiments on the kalantari et al.’s[18] dataset using the AHDR[49] baseline. Specifically, we integrated semantic information in two standard ways common in low-level vision. First is semantic feature injection: we extracted semantic features from each input SDR frame via a pre-trained FastSAM[64] encoder, fused them using an FPN, and used the resulting representation to modulate AHDR’s intermediate features through affine transformations. Second is semantic segmentation map guidance: we generated per-SDR segmentation masks with FastSAM and concatenated them with SDR inputs as additional channels to guide HDR reconstruction. As shown in Table IV, both variants yield a modest improvement in PSNR-PU but still lag significantly behind our proposed framework. We attribute this to two key challenges in multi-exposure HDR reconstruction: (1) Input SDR images suffer from exposure-related degradations, such as under/over-exposure and noise, which makes direct extraction of reliable semantic priors difficult. (2) Frame misalignment, caused by camera or object motion, leads to inconsistencies in per-frame semantic estimates and conflicting guidance during fusion. In contrast, our method extracts semantic priors from a static, initial HDR estimate. This estimate integrates information from all exposures, yielding a more robust and coherent semantic representation that is less affected by degradations or motion.
Figure 8: Segmentation results of HDR images from different domains and formats in FastSAM.
IV-E Discussion
Computational Overhead of Training Stage. In the proposed framework, we utilize semantic knowledge extracted from a pre-trained semantic model to tackle the ill-posed problem of HDR reconstruction. Upon completion of training, our approach substantially improves imaging quality without necessitating alterations to the original model architecture or introducing additional computational overhead during inference. However, the training phase incurs extra computational costs due to the need to train the Semantic Prior Guided Reconstruction Module and conduct inference using FastSAM. We detail the training computational overhead in Table V, based on an NVIDIA A100 GPU with 128×128 image patches and a batch size of 8. The primary source of this overhead is the SPGRM, which requires preprocessing of semantic prior features and reconstructs high-quality HDR images under their guidance. In comparison, the Semantic Knowledge Alignment Module imposes a relatively minor computational burden, as it conducts feature semantic alignment in the latent space using only a small number of convolutional layers. Additionally, while FastSAM possesses a significantly larger parameter count than the original model (68M vs. 1.24M), its weights are frozen during training, eliminating the need for gradient computations and restricting its role to inference. Consequently, FastSAM contributes minimal computational overhead and negligible memory usage. When using AHDR as the baseline model, the total time to complete the entire training process on a single NVIDIA A100 GPU is approximately 20 hours. This total duration varies depending on the chosen baseline model. For instance, if the baseline model is inherently computationally intensive (e.g., HDR-Transformer [31]), the relative proportion of the additional overhead introduced by our framework in the total time will be comparatively smaller.
Overall, although our framework introduces increased computational demands and memory consumption during training stage, it offers a notable advantage in resource efficiency compared to retraining a semantic prior model tailored specifically for HDR tasks. Furthermore, in the field of HDR reconstruction, existing methods under the same paradigm typically improve model performance at the cost of additional inference overhead. In contrast, our approach achieves comparable improvements without modifying the network architecture, thereby avoiding any extra computational overhead during inference.
Choice of Domain Transfer Operation. Due to the inherent differences in domain and format between HDR and SDR images, our framework requires preprocessing HDR images to obtain priors and aligning the outputs of different models before performing distillation. The choice of domain transfer operation for this alignment is critical to the success of our approach. For sRGB HDR images, this involves a tone mapping operation, while for RAW images, an additional demosaicing operation is included.
In our method, we adopt the $\mu$-law for tone mapping due to its proven effectiveness and simplicity. The $\mu$-law function has been widely utilized in prior work [49, 17, 34] for loss function computations, demonstrating its robustness. Although more advanced tone mapping techniques exist, they are often complex and non-differentiable, making them unsuitable for our framework. For instance, Gamma encoding, despite its simplicity, is non-differentiable near zero, limiting its applicability. Additionally, converting RAW images to sRGB requires multiple operations, such as demosaicing and white balancing. Among these, demosaicing is particularly crucial, as it directly impacts the structural integrity of the image. In contrast, white balancing relies on additional metadata and primarily affects the image’s color tone without altering its content. Our experiments show that white balancing has minimal impact on segmentation results. Consequently, we incorporate only an additional demosaicing operation alongside tone mapping to process RAW images effectively. As illustrated in Fig. 8, HDR images processed with the proposed domain transfer operation can be effectively utilized for semantic knowledge extraction, demonstrating the robustness of our method.
Discussion on GAN-based Baselines. While our framework demonstrated robust improvements on CNN-based (e.g., AHDR [49]) and Transformer-based (e.g., HDR-Trans [31]) models, we encountered challenges when applying it to GAN-based methods, specifically HDRGAN [35], which is widely recognized as an excellent and influential method in the field of HDR reconstruction. Despite its theoretical merits, we observed severe training instability during our integration experiments. This sensitivity is likely due to the inherent difficulty in balancing the generator and discriminator during adversarial training, which requires extensive hyperparameter tuning and careful optimization scheduling. We report this negative result to highlight the potential limitations of applying our current framework to GAN-based models, suggesting that specialized stabilization techniques may be required for such architectures.
Limitations and Future Work A key limitation of the proposed framework lies in the increased computational overhead during training: compared to the original reconstruction model, the SPGRM and semantic feature preprocessing add extra training time, though inference remains efficient by only using the optimized ORM. Another limitation is its limited support for HDR-specific wide-gamut color spaces (e.g., Rec. 2020, DCI-P3) that are widely used in professional HDR scenarios, current experiments focus on sRGB/RAW formats, and the existing $\mu$-law domain transfer cannot fully leverage the dynamic range of these advanced color spaces. For future work, we will explore more lightweight knowledge transfer strategies to reduce training overhead while preserving semantic cues, and extend the domain transfer module to support Rec. 2020/DCI-P3, further validating the framework on professional HDR datasets.
TABLE V: Additional computational overhead of training stage.
Module Para. Flops GPU Memory Time
AHDR 1.24M 163.45G 2.38GB 31.8ms
SPGRM+FPN 1.44M 151.09G 4.76GB 69.5ms
SKAM 0.67M 87.32G 0.88GB 17.6ms
FastSAM 68M 76.77G 0.32GB 30.19ms
V Conclusion
@@
Existing multi-exposure HDR imaging methods primarily focus on mitigating artifact issues caused by motion, but they overlook the inherently ill-posed nature of captured SDR images in real-world scenarios. To tackle these challenges, we introduce a general framework that leverages semantic priors from pretrained segmentation models to boost existing HDR reconstruction models without modifying the original network. By endowing models with semantic knowledge, our framework enables a deeper understanding of image content, providing explicit guidance for the reconstruction process. Extensive experiments demonstrate that the proposed framework consistently improves the performance of multiple baseline models, achieving state-of-the-art results across diverse datasets encompassing various scenes and formats.
Acknowledgment
This work is supported by NSFC of China under Grant 62301432 and Grant 62306240, the Natural Science Basic Research Program of Shaanxi under Grant 2023-JC-QN-0685 and Grant QCYRCXM-2023-057, the Fundamental Research Funds for Central Universities, and Guangdong Basic and Applied Basic Research Foundation 2025A1515011119, the Innovation Foundation for Doctor Dissertation of Northwestern Polytechnical University ZX2025019.